OpenStudy (anonymous):

f(x) = Square root of quantity x plus nine.; g(x) = 8x - 13 Find f(g(x))

OpenStudy (anonymous):

\[f(x) = x^2 + 9\] \[g(x) = 8x - 13\] To find f(g(x)) replace x of f(x) with 8x - 13.. That is: \[f(8x - 13) = (8x - 13)^2 + 9\] Solve it by using the formula: \[(a - b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

thank you this helped a lot

OpenStudy (anonymous):

Welcome dear..